OpenStudy (anonymous):

You own 6 cassettes and are taking 5 on vacation. In how many ways can you choose 5 cassettes from the 6?

OpenStudy (lgbasallote):

use combinations..i guess \[nCr = \frac{n!}{r!(n-r)!}\] so \[6C5 = \frac{6!}{5!(6-5)!}\]

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

will u explain it a bit further

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (lgbasallote):

which do you need explaining?

OpenStudy (anonymous):

what do udo after thaat

OpenStudy (anonymous):

will u explain it a bit further

OpenStudy (anonymous):

what do udo after thaat

OpenStudy (lgbasallote):

you solve it...you do know what the "!" means right?

OpenStudy (anonymous):

yes its like 6x5x4x3x2x1